# Scaling laws in wearable human activity recognition

### v1

@@ Abstract
Many deep architectures and self-supervised pre-training techniques have been proposed for human activity recognition (HAR) from wearable multimodal sensors. Scaling laws have the potential to help move towards more principled design by linking model capacity with pre-training data volume. Yet, scaling laws have not been established for HAR to the same extent as in language and vision. By conducting an exhaustive grid search on both amount of pre-training data and Transformer architectures, we establish the first known scaling laws for HAR. We show that pre-training loss scales with a power law relationship to amount of data and parameter count and that increasing the number of users in a dataset results in a steeper improvement in performance than increasing data per user, indicating that diversity of pre-training data is important, which contrasts to some previously reported findings in self-supervised HAR. We show that these scaling laws translate to downstream performance improvements on three HAR benchmark datasets of postures, modes of locomotion and activities of daily living: UCI HAR and WISDM Phone and WISDM Watch. Finally, we suggest some previously published works should be revisited in light of these scaling laws with more adequate model capacities.
@@ 1 Introduction
RQ0: Does the Vision Transformer (ViT) architecture generalize seamlessly to the time series domain of wearable motion sensing?
@@
RQ2: How does data diversity affect the scaling laws and how does it align with previously published work?
@@
  • We summarize the state of the art in HAR, including self-supervised training, and in the work to uncover scaling laws in foundation models. We identify that scaling laws have not yet been investigated in HAR (Section 2).
  • We introduce a ViT architecture adapted for HAR from triaxial accelerometer and gyroscope (6 channel input) which is trained with self-supervised principles. We faithfully apply Masked Autoencoder pre-training and ViT encoder architecture with minimal changes from the original vision papers, addressing RQ0. For our scaling laws, we differ from language and vision domains by training to convergence in order to study the relationship between number of parameters and data more directly under the unique conditions for HAR, where we are not constrained by training compute (Section 3).
@@
  • We indicate how these scaling laws can be used to inform future research. Notably, we discuss how these scaling laws can be used to re-interpret previously published work, addressing RQ3. We give two examples where we assess that increased model capacity could have been beneficial (Section 5). Our findings reiterate the importance of pre-training with higher “quality” data rather than solely larger volume of data: we show that adding more users to the pre-training dataset is better than adding more data from the same user, contradicting other recent findings in the field (e.g., Narayanswamy et al. (2024)), and pointing to future research avenues, addressing RQ2.
@@ 2.2 Self-supervised learning for HAR
Annotation scarcity in activity recognition can be combated through self-supervised learning, where a pretext task is learnt on an unannotated dataset, with fine-tuning on a smaller annotated dataset. An exhaustive review of methods can be found in Haresamudram et al. (2022); Logacjov (2024). These reviews highlight that variations of masked reconstruction are commonly used. Haresamudram et al. (2020) proposes a Transformer encoder in the time domain, yielding a model with 1.5M parameters. They only mask 10% of data and use an MLP decoder. SelfPAB Logacjov et al. (2024) Logacjov and Bach (2024) and FreqMAE Kara et al. (2024) use masked autoencoder pre-trained on the spectrograms of the input. SelfPAB is inspired by audio models, and does pre-training on the HUNT4 dataset with 100k hours of sensor signals. FreqMAE is suggesting a specific Transformer block to account for spectral properties of the input signal. Our approach differs from these by faithfully applying masked autoencoder He et al. (2022) with minimal change from the vision domain, demonstrating that this approach generalizes well to the HAR domain.
@@ 2.3 Scaling laws in language, vision and HAR
Scaling laws have been explored in sensor foundation models Narayanswamy et al. (2024); Abbaspourazad et al. (2024). The work in Narayanswamy et al. (2024) differs fundamentally from ours. The authors create a foundational model not on raw motion sensor data but on a set of 10 engineered statistical features extracted from the motion sensors at a rate of one vector every minute, as well as physiological sensors. Although they introduce scaling laws, operating on engineered features makes it difficult to draw direct parallels to our work. Furthermore, they rely on a proprietary dataset, and report scaling in terms of reconstruction performance (mean squared error) instead of downstream classification performance. Our work instead uses public datasets to help reproducibility, and our conclusions differ from theirs regarding data diversity (Section 5). Abbaspourazad et al. (2024) mention scaling up models as a future research direction and do not yet draw scaling laws from their experiments.
@@ 3.1 Scaling Laws
For our scaling laws, we take a different approach from that in language Hoffmann et al. (2022) Kaplan et al. (2020) and vision Zhai et al. (2022). In these domains, since data was abundant and compute was the primary constraint, they never completed a full epoch, and thus equated number of steps to amount of data. For HAR, however, data is the primary constraint, so we repeat data many times (over 100 epochs) until convergence. Since we are able to train even our largest models to convergence, we do not fix the amount of compute. Instead, we focus on the capacity of the models (number of parameters). This is more directly tied to inference cost than training, which aligns with the priorities of many HAR deployments.
@@ 3.5 Datasets
For downstream evaluation, we use popular benchmark datasets UCI HAR and WISDM Phone/Watch datasets. We use the full training dataset for all supervised training. UCI HAR Reyes-Ortiz and Parra (2013) contains data from 30 volunteers aged 19-48 engaging in 6 modes of locomotion and postures: walking, walking upstairs, walking downstairs, sitting, standing, laying. The accelerometer and gyroscope data is recorded at 50Hz from a smartphone worn on the waist. We use the same random partitioning prescribed by the dataset authors (70% training and 30% test sets). We also keep the existing raw data preprocessing pipeline, involving noise filtering, 2.56sec sliding windows with 50% overlap, resulting in 128 samples per window, and use the raw acceleration, not the low-pass filtered version also present in the dataset.
@@ 4.3 Downstream performance
We show that the scaling laws for pre-training translate to similar trends in improved downstream linear classification performance. For each pre-training dataset size, we plot the best F1 score from linear evaluation on downstream datasets UCI HAR and WISDM Phone/Watch. This can be seen in Figure 4. Contrary to previous findings Haresamudram et al. (2022); Dhekane et al. (2023), we see consistent improvement as we scale the data size. For UCI HAR, we reach 97.9% F1 score with linear evaluation. To our knowledge, this is on par with the best reported result (98.6% from Nguyen et al. (2024)) for this dataset. For all datasets, we surpass from scratch baseline results, with significant improvement for phone datasets UCI HAR (+2.8pp) and WISDM Phone (+2.4pp). For WISDM Watch, the improvement is smaller (+0.5pp). This is not surprising given that our pre-training dataset consists of only phone data. Still, the consistent increase in watch performance suggests that we are seeing positive transfer between body positions.
@@ 4.5 Augmentations
We apply augmentations during pre-training, and study the effect on downstream model performance. In Figure 7 We separate results by encoder as in Figure 5, but take the best score both with and without augmentations. We see that augmentations always improve performance, especially at larger scales. The optimal capacity without augmentations can be smaller, at either the 4th largest model (about 10M parameters) for UCI HAR or the 3rd largest model (20M parameters) for WISDM Phone. This is not surprising, since augmentations can be thought of as a strong regularizer, and/or an artificial expansion of the dataset.
@@ 5 Conclusion
We demonstrate that Masked Autoencoder and ViT approaches generalize directly to the domain of HAR without the need for domain-specific design choices. This is beneficial to take advantage of the findings resulting from advances in ViT in other domains, and is unlike what is often pursued in wearable and mobile computing which tend to propose ad-hoc architectures.
@@
This contradicts the findings from Narayanswamy et al. (2024) regarding increasing the diversity of data: our findings show this to be beneficial, while theirs suggest increasing volume without necessarily increasing diversity. This may be due to differences in their evaluations, and indicates the need for further investigation. Indeed, we operated on raw sensor data (instead of their model operating on top of 10 statistical features extracted from raw data) and we evaluated performance on downstream task with a much larger number of activity classes (18 compared to 8 in their work).

### v2

@@ Abstract
Many deep architectures and self-supervised pre-training techniques have been proposed for human activity recognition (HAR) from wearable multimodal sensors. Scaling laws have the potential to help move towards more principled design by linking model capacity with pre-training data volume. Yet, scaling laws have not been established for HAR to the same extent as in language and vision. By conducting an exhaustive grid search on both amount of pre-training data and Transformer architectures, we establish the first known scaling laws for HAR. We show that pre-training loss scales with a power law relationship to amount of data and parameter count and that increasing the number of users in a dataset results in a steeper improvement in performance than increasing data per user, indicating that diversity of pre-training data is important. We show that these scaling laws translate to downstream performance improvements on three HAR benchmark datasets of postures, modes of locomotion and activities of daily living: UCI HAR and WISDM Phone and WISDM Watch. Finally, we suggest some previously published works should be revisited in light of these scaling laws with more adequate model capacities.
@@ 1 Introduction
RQ2: How does data diversity affect the scaling laws?
@@
  • We indicate how these scaling laws can be used to inform future research. Notably, we discuss how these scaling laws can be used to re-interpret previously published work, addressing RQ3. We give two examples where we assess that increased model capacity could have been beneficial (Section 5). Our findings reiterate the importance of pre-training with greater “diversity” data rather than solely larger volume of data: we show that adding more users to the pre-training dataset is better than adding more data from the same user.
@@ 2.2 Self-supervised learning for HAR
Annotation scarcity in activity recognition can be combated through self-supervised learning, where a pretext task is learnt on an unannotated dataset, with fine-tuning on a smaller annotated dataset. An exhaustive review of methods can be found in Haresamudram et al. (2022); Logacjov (2024). These reviews highlight that variations of masked reconstruction are commonly used. Haresamudram et al. (2020) proposes a Transformer encoder in the time domain, yielding a model with 1.5M parameters. They only mask 10% of data and use an MLP decoder. SelfPAB Logacjov et al. (2024) Logacjov and Bach (2024) and FreqMAE Kara et al. (2024) use masked autoencoder pre-trained on the spectrograms of the input. SelfPAB is inspired by audio models, and does pre-training on the HUNT4 dataset with 100k hours of sensor signals. FreqMAE is suggesting a specific Transformer block to account for spectral properties of the input signal. Our approach differs from these by faithfully applying masked autoencoder He et al. (2022) with minimal change from the vision domain.
@@ 2.3 Scaling laws in language, vision and HAR
Scaling laws have been explored in sensor foundation models Narayanswamy et al. (2024); Abbaspourazad et al. (2024). The work in Narayanswamy et al. (2024) differs fundamentally from ours. The authors create a foundational model not on raw motion sensor data but on a set of 10 engineered statistical features extracted from the motion sensors at a rate of one vector every minute, as well as physiological sensors. Although they introduce scaling laws, operating on engineered features makes it difficult to draw direct parallels to our work. Furthermore, they rely on a proprietary dataset, and report scaling in terms of reconstruction performance (mean squared error) instead of downstream classification performance. Our work instead uses public datasets to help reproducibility, and we introduce new conclusions regarding data diversity (Section 5). Abbaspourazad et al. (2024) mention scaling up models as a future research direction and do not yet draw scaling laws from their experiments.
@@ 3 Method
In this paper we explore scaling laws based on Masked Autoencoder with a ViT encoder for the following reasons:
  • In language and vision domains, Transformers and Masked Autoencoder have proven to scale well due to high paralellization (e.g. compared to recurrent nets). Furthermore, scaling laws in language and vision were studied using similar architectures, which enables us to draw parallels between domains.
  • Narayanswamy et al. (2024) used this architecture on coarse statistical engineered features. We aim to study scaling laws on the raw sensor signal.
  • This architecture has been thoroughly vetted for HAR and is in wide spread use Haresamudram et al. (2020) Kara et al. (2024). By studying scaling laws on this architecture we ensure broader applicability of our findings Section 5.
@@ 3.1 Scaling Laws
For our scaling laws, we take a different approach from that in language Hoffmann et al. (2022) Kaplan et al. (2020) and vision Zhai et al. (2022). In these domains, since data was abundant and compute was the primary constraint, they never completed a full epoch, and thus equated number of steps to amount of data. For HAR, however, data is the primary constraint, so we repeat data many times (over 100 epochs) until convergence. We fix the number of pre-training steps to 500,000, which was found experimentally to be sufficient for convergence given our model and data sizes. Since we are able to train even our largest models to convergence, we do not fix the amount of compute. Instead, we focus on the capacity of the models (number of parameters). This is more directly tied to inference cost than training, which aligns with the priorities of many HAR deployments.
@@ 3.5 Datasets
For downstream evaluation, we use popular benchmark datasets UCI HAR and WISDM Phone/Watch datasets. We use the full training dataset for all supervised training. UCI HAR Reyes-Ortiz et al. (2013) contains data from 30 volunteers aged 19-48 engaging in 6 modes of locomotion and postures: walking, walking upstairs, walking downstairs, sitting, standing, laying. The accelerometer and gyroscope data is recorded at 50Hz from a smartphone worn on the waist. We use the same random partitioning prescribed by the dataset authors (70% training and 30% test sets). We also keep the existing raw data preprocessing pipeline, involving noise filtering, 2.56sec sliding windows with 50% overlap, resulting in 128 samples per window, and use the raw acceleration, not the low-pass filtered version also present in the dataset.
@@ 4.3 Downstream performance
We show that the scaling laws for pre-training translate to similar trends in improved downstream linear classification performance. For each pre-training dataset size, we plot the best F1 score from linear evaluation on downstream datasets UCI HAR and WISDM Phone/Watch. This can be seen in Figure 4. Contrary to previous findings Haresamudram et al. (2022), we see consistent improvement as we scale the data size. For UCI HAR, we reach 97.9% F1 score with linear evaluation. To our knowledge, this is on par with the best reported result (98.6% from Nguyen et al. (2024)) for this dataset. Note that we do not compare with any public WISDM results because this dataset lacks a common protocol (e.g. train/test split), making direct comparisons irrelevant. For all datasets, we surpass from scratch baseline results, with significant improvement for phone datasets UCI HAR (+2.8pp) and WISDM Phone (+2.4pp). Note that WISDM Phone results are low overall, but this can be explained by our protocol including all classes, some of which are difficult or impossible to detect from phone alone (e.g. typing). For WISDM Watch, the improvement is smaller (+0.5pp). This is not surprising given that our pre-training dataset consists of only phone data. Still, the consistent increase in watch performance suggests that we are seeing positive transfer between body positions.
@@ 4.5 Augmentations
We apply random rotation and scaling augmentations during pre-training, and study the effect on downstream model performance. In Figure 7 We separate results by encoder as in Figure 5, but take the best score both with and without augmentations. We see that augmentations always improve performance, especially at larger scales. The optimal capacity without augmentations can be smaller, at either the 4th largest model (about 10M parameters) for UCI HAR or the 3rd largest model (20M parameters) for WISDM Phone. This is not surprising, since augmentations can be thought of as a strong regularizer, and/or an artificial expansion of the dataset.